# Implicit Neural Representations for Spatial Transcriptomics

### v1

@@ Abstract
Spatial Transcriptomics (ST) is a method that captures spatial gene expression profiles within histological sections. The discrete spatial distribution and the super-high dimensional sequencing results make ST data challenging to be modeled effectively. In this paper, we manage to model ST in a continuous and compact manner by the proposed tool, SUICA, empowered by the great approximation capability of Implicit Neural Representations (INRs) that can improve both the spatial resolution and the gene expression. Concretely within the proposed SUICA, we incorporate a graph-augmented Autoencoder to effectively model the context information of the unstructured spots and provide informative embeddings that are structure-aware for spatial mapping. We also tackle the extremely skewed distribution in a regression-by-classification fashion and enforce classification-based loss functions for the optimization of SUICA. By extensive experiments of a wide range of common ST platforms, SUICA outperforms both conventional INR variants and SOTA methods for ST super-resolution regarding numerical fidelity, statistical correlation, and bio-conservation. The prediction by SUICA also showcases amplified gene signatures that enriches the bio-conservation of the raw data and benefits subsequent analysis. The code is available at https://github.com/Szym29/SUICA.
@@ 1 Introduction
Spatial Transcriptomics (ST) enables scientists to quantify gene expression while preserving spatial information in tissue sections [27]. These platforms use various strategies to capture mRNA transcripts from tissue sections and perform sequencing to quantify the gene expressions at spatially defined locations [31]. However, achieving higher resolution is essential for accurately modeling and analyzing cellular functions [54], while for now, ST data can still be rather expensive (\$3,500/$cm^{2}$ for the capture chip and \$800/$cm^{2}$ for the high-density sequencing [2]). Beyond the cost constraints to get high-resolution ST, another key technical challenge lies in the drop-out rates (high sparsity of ST data) that makes the gene expressions less informative for analysis [42]. Compounding this challenge, different sequencing techniques and platforms exhibit significant heterogeneity in terms of spatial distribution, sequencing depth (number of mRNA readouts per spot), and drop-out rates [31]. As a result, such the difference makes a general backbone to model and analyze data across different tissues by different platforms less feasible. To address these technical challenges, we propose a computational framework that transforms discrete ST data into a continuous, compact representation. Our approach enables super-resolution spatial transcriptomics analysis, providing a robust foundation for downstream analyses.
Figure 1: Starting with the discretely sampled spots (a) of ST, SUICA performs continuous modeling (b) by aid of the great approximation power of INRs. This approach enables more precise profiling of cell heterogeneity, as visualized in the UMAP (c), which further facilitating the discovery of new biology.
@@
Recently, Implicit Neural Representations (INRs) have drawn great attention from researchers for their compact, continuous, and differentiable properties as a novel representation of general coordinate-based signals. INRs map coordinates to corresponding values with a neural network with wide applications in inverse graphics [29], geometric modeling [38], and video compression [4]. Motivated by INRs’ inherent smoothness property, we leverage this characteristic to interpolate between sparse ST spots, enabling more comprehensive spatial transcriptomics analysis. However, applying INRs to ST data faces two major challenges. First, scaling INRs to the super-high dimensional space of gene expression is non-trivial, as simply widening or deepening the network cannot effectively overcome the curse of dimensionality. Second, the low mRNA transcript capture rate in ST, combined with varying gene expression patterns across cellular states, results in zero-inflated ST data with high sparsity [41]. This sparsity makes it particularly challenging for conventional INRs to accurately capture the underlying complex, non-linear spatial patterns.
@@
Building upon this approach, we introduce SUICA, a powerful variant of INR designed especially for ST data. SUICA fully accounts for the unique properties of ST and the diversity of current mainstream ST platforms. In SUICA, we employ an Autoencoder (AE) based on Graph Convolutional Network (GCN) [18] to bridge the gap between conventional INRs and ST. The GCN enhances the approximation capability for unstructured data by leveraging contextual information and making the embeddings structure-aware. By leveraging the strong profiling power of Graph Autoencoder (GAE) regarding the sparse and skewed distribution, we perform the fitting of INRs at the expressive low-dimensional embedding space, which has been proven to be compact and more suitable for INRs. This approach facilitates more accurate and efficient modeling of ST data.
@@
To optimize SUICA and enforce the sparsity within the regression-based scheme of INRs, we construct pseudo-probabilities and adopt a regression-by-classification approach for training. With SUICA, we are enabled to model diverse ST data from several popular platforms and sequencing protocols using a unified representation, facilitating comprehensive analysis as illustrated in Fig. 1. Extensive experimental results demonstrate that SUICA outperforms existing INR variants as well as the state-of-the-art methods in both numerical fidelity and statistical correlation. Moreover, SUICA’s imputation ability leads to more cell-type informed clustering results of ST. To summarize, our contributions in this paper are three-fold as follows:
@@
  • We introduce SUICA, modeling ST data as a continuous and compact representation while preserving data authenticity;
@@ 2.1 Implicit Neural Representations
The idea of neural networks for function approximation dates back to the 1980s and 1990s [6, 14] as universal function approximation theories. From a relatively modern perspective, SIREN [43] adopts periodic sine functions as the activation function and learns INRs rich in fine details. Equipped with Random Fourier Features, FFN [48] multiplies the coordinates with a random feature sampled from a normal distribution. BACON [23] further explores the spectral bandwidth of the target signal and progressively regresses in a coarse-to-fine manner.
INRs have been applied across various downstream tasks, including inverse graphics and video compression. NeRF [29] parameterizes the 3D scene as an MLP-based radiance field with colors and densities. It applies differential volume rendering to query a series of points in the field and then calculate the weighted sum as the pixel value. Following this parameterization, subsequent works [24, 56, 46, 16, 11, 32, 40, 1, 3] further improve the radiance field for efficient training and rendering. Additionally, the continuity of INR is well-suited for application in 3D geometric reconstruction methods [38, 52, 55, 33]. As for video compression, NeRV [4] and subsequent variants [22, 5] employ compact INRs to reduce data redundancy.
It is noteworthy that existing INR applications map from low-dimensional to low-dimensional spaces, e.g., $\mathbb{R}^{2}\to\mathbb{R}^{3}$ for image regression, and $\mathbb{R}^{5}\to\mathbb{R}^{4}$ for inverse graphics. The use of INRs for mapping from low dimension (i.e., $\mathbb{R}^{2}$) to super-high dimension (i.e., over 20,000 channels) remains unexplored.
Figure 2: The overall pipeline of SUICA. SUICA is a powerful tool powered by INRs to improve both spatial resolution and cell profiles (gene expression).
@@ 2.2 Deep Learning for Spatial Transcriptomics
Deep learning methods are increasingly adopted in ST to enhance spatial resolution, denoise raw data, improve spatial domain clustering, and support more accurate downstream analysis [57]. Given the complexity of capturing cell-to-cell and spatial dependencies, graph-based methods leveraging cell neighborhood relationships have gained significant traction. Notable examples include GCN-based methods like SpaGCN [15], STAGATE [9], and GraphST [25], as well as graph transformer-based model like SiGra [49]. These graph-based methods progressively refine the modeling of ST data, enabling more nuanced interpretation of complex tissue structure.
ST data often require a higher resolution to accurately capture fine spatial patterns within tissues. Several methods leverage histological images to enhance the spatial resolution of gene expression data, including Convolution Neural Network (CNN)-based models ST-Net [12] and DeepSpaCE [30], and Vision Transformer-based methods such as HistoGene [37] and Hist2ST [58]. In cases without histological images, other approaches  [59, 61, 21] rely solely on spatial and gene expression data to improve resolution. To generate unmeasured spots profiles, STAGE [21] trains a spatial location-supervised Autoencoder that enhances resolution by integrating spatial coordinates with gene expression data. In summary, these resolution enhancement methods significantly improve ST quality, improving spatial visualization and facilitating more accurate downstream analyses.
@@ 3.1.1 Spatial Transcriptomics
ST data can be viewed as a point cloud, where each point is parameterized by spatial coordinates $\mathbf{x}$. The mRNA readouts are bounded at these sampled locations and represented by a high dimensional vector, where each channel reflects the numerical abundance of a particular mRNA transcript (0 indicates absence or failed capture).
In terms of the spatial distribution of the sampled spots, different ST platforms follow different sampling protocols, e.g., Visium applies a fixed array of probes so the sampled points are distributed uniformly over the whole slice, whereas other common platforms do not guarantee such regularity. Thus, to accommodate general use cases, ST data is often modeled as unstructured without quantizing into regular grids, making CNN-based analysis inapplicable.
@@ 3.1.2 Challenges
The key challenge in modeling ST lies in the super-high dimensional representation, often exceeding 20,000 channels. Due to low mRNA capture rates and varying gene expression patterns across cellular states, ST data is typically highly sparse, with zero values comprising up to 90%. A satisfying model must maintain the inherent sparsity of ST data, as it reflects the true gene expression. However, vanilla INRs are observed to have the tendency to yield normally distributed outputs that are smooth [20], rather than the zero-inflated ones of ST, presenting a significant challenge.
Beyond maintaining sparsity, it is also crucial to ensure the numerical fidelity of non-zero values for accurate cell type identification. It is worth noting that high sparsity complicates evaluation of reconstruction quality, as an entirely empty prediction may still result in low loss, yet be entirely unacceptable.
To this end, the design of the representations and evaluation protocol must address both sparsity and numerical fidelity—an aspect that has, to our knowledge, not been fully explored in the context of INRs.
@@ 3.2.1 Overview
The overall pipeline is illustrated in Fig. 2. To construct a compact and dense embedding domain for subsequent INRs, we first incorporate a graph-based encoder and pre-train Graph Autoencoder (GAE) using the given ST slice in a self-regressing manner. With the pre-trained GAE, we obtain the encoded latent representation for all spots, denoted as $\mathbf{z}_{\text{gt}}$.
We then initiate INR tuning to optimize the neural mapping from ST coordinates to the embeddings, targeting $\mathbf{z}_{\text{gt}}$. Once the INR reaches a stable state, we attach a decoder that is pre-trained in our GAE. Subsequently, we fix the INR and train the decoder to fit the raw dataset readouts, $\mathbf{y}_{\text{gt}}$.
@@ 3.2.2 Graph-augmented Autoencoder
The sequenced readouts of ST are known for an extremely skewed distribution, represented as high-dimensional, sparse data, making them challenging for vanilla INRs to model effectively. This sparsity exacerbates the curse of dimensionality: as dimensionality increases, the volume of the data space expands rapidly, leading to sparsely distributed data. Consequently, the data required for reliable results grows exponentially with dimensionality, rendering data points increasingly dissimilar and challenging to organize efficiently.
@@
As illustrated in Fig. 3, we conduct a graph-based spectral analysis of embeddings generated by both the AE and the GAE. Specifically, we construct a connectivity graph using Euclidean distance between spots ($k=4$) and calculate the Graph Total Variation (GTV) for embeddings defined on the built graph. Fig. 3 is colored based on aggregated edge-wise variations of each vertex, revealing that the GAE produces shaper structural clues. Additionally, we compute the channel-wise variance for both embeddings. Together with the GTV results, these findings demonstrate that GAE offers a more expressive and informative representation, effectively capturing disentangled embeddings for ST.
@@
where $\mathbf{M}_{\mathbf{y}}$ represents the binary mask over $\mathbf{y}$.
@@ 3.2.3 Embedding Mapping
For SUICA, we empirically select two baseline architectures: FFN [48] and SIREN [43], both of which have demonstrated effectiveness across various data types. In our experiments, we apply SIREN for spatially sparse ST and FFN for denser spatial distributions. A detailed discussion on this selection is included in the supplementary material.
Following conventional INR approaches, we assume Gaussian-distributed error in the predicted $\hat{\mathbf{z}}$ and use MSE loss for optimization:
$\mathcal{L}_{\text{embd}}=\frac{1}{|\mathbf{M}_{\mathbf{z}}|}\sum_{\mathbf{M}_{\mathbf{z}}}(\hat{\mathbf{z}}-\mathbf{z}_{\text{gt}})^{2},$ (2)
where $\mathbf{M}_{\mathbf{z}}$ represents the binary mask over elements in $\mathbf{z}$. While $\hat{\mathbf{z}}$ can serve as the mapping output for downstream analysis, we consider it as an intermediate result, decoding this latent code back into the raw space to achieve a more comprehensive reconstruction.

### v2

@@ Abstract
Spatial Transcriptomics (ST) is a method that captures gene expression profiles aligned with spatial coordinates. The discrete spatial distribution and the super-high dimensional sequencing results make ST data challenging to be modeled effectively. In this paper, we manage to model ST in a continuous and compact manner by the proposed tool, SUICA, empowered by the great approximation capability of Implicit Neural Representations (INRs) that can enhance both the spatial density and the gene expression. Concretely within the proposed SUICA, we incorporate a graph-augmented Autoencoder to effectively model the context information of the unstructured spots and provide informative embeddings that are structure-aware for spatial mapping. We also tackle the extremely skewed distribution in a regression-by-classification fashion and enforce classification-based loss functions for the optimization of SUICA. By extensive experiments of a wide range of common ST platforms under varying degradations, SUICA outperforms both conventional INR variants and SOTA methods regarding numerical fidelity, statistical correlation, and bio-conservation. The prediction by SUICA also showcases amplified gene signatures that enriches the bio-conservation of the raw data and benefits subsequent analysis. The code is available at https://github.com/Szym29/SUICA.
Spatial Transcriptomics, Implicit Neural Representations
Figure 1: Starting with the discretely sampled spots (a) of ST, SUICA performs continuous modeling (b) by aid of the great approximation power of INRs. This approach enables complete profiling of cell heterogeneity, as visualized in the UMAP (c), which further facilitating the discovery of new biology.
@@ 1 Introduction
Spatial Transcriptomics (ST) enables scientists to quantify gene expression while preserving spatial information in tissue sections (Marx, 2021). These platforms use various strategies to capture mRNA transcripts from tissue sections and perform sequencing to quantify the gene expressions at spatially defined locations (Moses & Pachter, 2022). However, various degradation patterns are being observed due to some limitations throughout the data acquisition process. For instance, achieving higher resolution is essential for accurately modeling and analyzing cellular functions (Williams et al., 2022), while for now, ST data can still be rather expensive, e.g., \$3,500/$cm^{2}$ for the capture chip and \$800/$cm^{2}$ for the high-density sequencing according to Chen et al. (2022a). Beyond the cost constraints to get high-resolution ST, another key technical challenge lies in the drop-out rates (high sparsity of ST data) that makes the gene expressions less informative for analysis (Qiu, 2020). Compounding this challenge, different sequencing techniques and platforms exhibit significant heterogeneity in terms of spatial distribution, sequencing depth (number of mRNA readouts per spot), and drop-out rates (Moses & Pachter, 2022). As a result, such the difference makes a general backbone to model and analyze data across different tissues by different platforms less feasible. To address these technical challenges, we propose a computational framework that transforms discrete ST data into a continuous, compact representation, and thus enables complete reconstruction of ST slides by recovering the degradations in the raw data without the need to know the specific type of degradation.
@@
Recently, Implicit Neural Representations (INRs) have drawn great attention from researchers for their compact, continuous, and differentiable properties as a novel representation of general coordinate-based signals. INRs map coordinates to corresponding values with a neural network with wide applications in inverse graphics (Mildenhall et al., 2020), geometric modeling (Park et al., 2019), and video compression (Chen et al., 2021). Motivated by INRs’ inherent smoothness property, we leverage this characteristic to interpolate between ST spots, enabling more comprehensive spatial transcriptomics analysis. However, applying INRs to ST data faces two major challenges. First, scaling INRs to the super-high dimensional space of gene expression is non-trivial, as simply widening or deepening the network cannot effectively overcome the curse of dimensionality. Second, the low mRNA transcript capture rate in ST, combined with varying gene expression patterns across cellular states, results in zero-inflated ST data with high sparsity (Piwecka et al., 2023). This sparsity makes it particularly challenging for conventional INRs to accurately capture the underlying complex, non-linear spatial patterns.
@@
Building upon this approach, we introduce SUICA, a powerful variant of INR designed especially for ST data. SUICA fully accounts for the unique properties of ST and the diversity of current mainstream ST platforms. In SUICA, we employ an Autoencoder (AE) based on Graph Convolutional Network (GCN) (Kipf & Welling, 2017) to bridge the gap between conventional INRs and ST. The GCN enhances the approximation capability for unstructured data by leveraging contextual information and making the embeddings structure-aware. By the aid of the strong profiling power of Graph Autoencoder (GAE) regarding the sparse and skewed distribution, we perform the fitting of INRs at the expressive low-dimensional embedding space, which has been proven to be compact and more suitable for INRs. To enforce the sparsity within the regression-based scheme of INRs, we construct pseudo-probabilities and adopt a regression-by-classification approach for training.
@@
With SUICA, we are enabled to reconstruct diverse ST data from several popular platforms and sequencing protocols using a unified representation, facilitating comprehensive analysis as illustrated in Figure 1. In particular, we apply SUICA to enhance the spatial resolution (spatial imputation), to alleviate the drop-out rates (gene imputation), and to smooth the noisy expressions (denoising), respectively. Extensive experimental results demonstrate that SUICA outperforms existing INR variants as well as the state-of-the-art methods in both numerical fidelity and statistical correlation. Moreover, SUICA’s imputation ability leads to more cell-type informed clustering results of ST. To summarize, our contributions in this paper are three-fold as follows:
@@
  • We introduce SUICA to model ST data as a continuous and compact representation while preserving data authenticity;
@@ 2 Related Work
Figure 2: The overall pipeline of SUICA. At training-time, a GAE based on cell graphs is trained, with whose pre-trained decoder concatenated to a INR. The INR then maps spot coordinates to the corresponding gene expressions. SUICA is capable of performing spatial imputation, gene imputation, and denoising.
@@ 2.1 Implicit Neural Representations
The idea of neural networks for function approximation dates back to the 1980s and 1990s (Cybenko, 1989; Hornik, 1991) as universal function approximation theories. From a relatively modern perspective, SIREN (Sitzmann et al., 2020) adopts periodic sine functions as the activation function and learns INRs rich in fine details. Equipped with Random Fourier Features, FFN (Tancik et al., 2020) multiplies the coordinates with a random feature sampled from a normal distribution. BACON (Lindell et al., 2022) further explores the spectral bandwidth of the target signal and progressively regresses in a coarse-to-fine manner.
INRs have been applied across various downstream tasks, including inverse graphics and video compression. NeRF (Mildenhall et al., 2020) parameterizes the 3D scene as an MLP-based radiance field with colors and densities. It applies differential volume rendering to query a series of points in the field and then calculate the weighted sum as the pixel value. Following this parameterization, subsequent works (Liu et al., 2020; Yu et al., 2021; Sun et al., 2022; Hu et al., 2022; Fridovich-Keil et al., 2022; Müller et al., 2022; Peng et al., 2020; Chan et al., 2022; Chen et al., 2022b) further improve the radiance field for efficient training and rendering. Additionally, the continuity of INR is well-suited for application in 3D geometric reconstruction methods (Park et al., 2019; Wang et al., 2021; Yariv et al., 2021; Oechsle et al., 2021). As for video compression, NeRV (Chen et al., 2021) and subsequent variants (Li et al., 2022; Chen et al., 2023) employ compact INRs to reduce data redundancy. It is noteworthy that existing INR applications map from low-dimensional to low-dimensional spaces, e.g., $\mathbb{R}^{2}\to\mathbb{R}^{3}$ for image regression, and $\mathbb{R}^{5}\to\mathbb{R}^{4}$ for inverse graphics. The use of INRs for mapping from low dimension (i.e., $\mathbb{R}^{2}$) to super-high dimension (i.e., over 20,000 channels) remains unexplored.
@@ 2.2 Deep Learning for Spatial Transcriptomics
Deep learning methods are increasingly adopted in ST to enhance spatial resolution, denoise raw data, improve spatial domain clustering, and support more accurate downstream analysis (Zahedi et al., 2024). Given the complexity of capturing cell-to-cell and spatial dependencies, graph-based methods leveraging cell relationships have gained significant traction. Notable examples include GCN-based methods like SpaGCN (Hu et al., 2021), STAGATE (Dong & Zhang, 2022), GraphST (Long et al., 2023), and SEDR (Xu et al., 2024) as well as graph transformer-based model like SiGra (Tang et al., 2023). These graph-based methods progressively refine the modeling of ST data, enabling more nuanced interpretation of complex tissue structure.
ST data often require a higher resolution to accurately capture fine spatial patterns within tissues. Several methods leverage histological images to enhance the spatial resolution of gene expression data, including Convolution Neural Network (CNN)-based models ST-Net (He et al., 2020) and DeepSpaCE (Monjo et al., 2022), and Vision Transformer-based methods such as Hist2ST (Zeng et al., 2022) and TRIPLEX (Chung et al., 2024). In cases without histological images, other approaches  (Zhao et al., 2021, 2023; Li et al., 2024) rely solely on spatial and gene expression data to improve resolution. To generate unmeasured spots profiles, STAGE (Li et al., 2024) trains a spatial location-supervised Autoencoder that enhances resolution by integrating spatial coordinates with gene expression data. In summary, these resolution enhancement methods significantly improve ST quality, improving spatial visualization and facilitating more accurate downstream analyses.
@@ 3.1.1 Spatial Transcriptomics
ST data can be viewed as an unordered point set, where each point is parameterized by spatial coordinates $\mathbf{x}$. The mRNA readouts are bounded at these sampled locations and represented by a high dimensional vector, where each channel reflects the numerical abundance of a particular mRNA transcript (0 indicates absence or failed capture).
In terms of the spatial distribution of the sampled spots, different ST platforms follow different sampling protocols, e.g., Visium applies a fixed array of probes so the sampled points are distributed uniformly over the whole slice, whereas other common platforms do not guarantee such regularity. Thus, to accommodate general use cases, ST data is often modeled as unstructured without being quantized into regular grids to make CNN-based analysis inapplicable.
@@ 3.1.2 Challenges
The key challenge in modeling ST lies in the super-high dimensional representation, often exceeding 20,000 channels. Due to low mRNA capture rates and varying gene expression patterns across cellular states, ST data are typically highly sparse, with zero values comprising up to 90%. A satisfying model must maintain the inherent sparsity of ST data, as it reflects the true gene expression. However, vanilla INRs are observed to have the tendency to yield normally distributed outputs that are smooth (Lee et al., 2018), rather than the zero-inflated ones of ST, presenting a significant challenge. It is also crucial to ensure the numerical fidelity of non-zero values for accurate cell type identification. It is worth noting that high sparsity complicates the evaluation of reconstruction quality, as an entirely empty prediction may still result in low loss, yet be entirely unacceptable. To this end, the design of the representations and evaluation protocol must address both sparsity and numerical fidelity—an aspect that has, to our knowledge, not been fully explored in the context of INRs.
@@ 3.2.1 Overview
The overall pipeline is illustrated in Figure 2. To construct a compact and dense embedding domain for subsequent INRs, we first incorporate a graph-based encoder and pre-train Graph Autoencoder (GAE) using the given ST slice in a self-regressing manner. With the pre-trained GAE, we obtain the encoded latent representation for all spots, denoted as $\mathbf{z}_{\text{gt}}$. We then initiate INR tuning to optimize the neural mapping from ST coordinates to embeddings, targeting $\mathbf{z}_{\text{gt}}$. Once the INR reaches a stable state, we attach a decoder that is pre-trained in our GAE. Subsequently, we fix the INR and train the decoder to fit the raw dataset readouts $\mathbf{y}_{\text{gt}}$.
Concretely, we perform 3 different tasks with SUICA, namely spatial imputation, gene imputation, and denoising, whose data flows are summarized as follows.
  • Spatial Imputation: Say an ST slice $\{(\mathbf{x},\mathbf{y})\}$, whose data spots are split into a training subset and a test subset. With the training subset, we train SUICA (GAE+INR), and infer the test subset for evaluation.
  • Gene Imputation: We randomly mute a part of the gene expressions of the data matrix, fit SUICA with all of the data, and infer all of the $\mathbf{x}$.
  • Denoising: The data flow is basically the same as gene imputation, but with injected noise as the degradation.
@@ 3.2.2 Graph-augmented Autoencoder
The sequenced readouts of ST are known for an extremely skewed distribution, represented as high-dimensional, sparse data. This sparsity exacerbates the curse of dimensionality, rendering data points increasingly dissimilar and challenging to organize efficiently.
@@
As illustrated in Figure 3, we conduct a graph-based spectral analysis of embeddings generated by both the AE and the GAE. Specifically, we construct a connectivity graph using Euclidean distance between spots ($k=5$) and calculate the Graph Total Variation (GTV) for embeddings defined on the built graph. Figure 3 is colored based on aggregated edge-wise variations of each vertex, revealing that the GAE produces shaper structural clues. Additionally, we compute the channel-wise variance for both embeddings. Together with the GTV results, these findings demonstrate that GAE offers a more expressive and informative representation, effectively capturing disentangled embeddings for ST.
@@
where $\mathbf{M}_{\mathbf{y}}$ represents the mask over all of the elements in $\mathbf{y}$ to compute the element-wise average.
@@ 3.2.3 Embedding Mapping
For SUICA, we empirically select two baseline architectures: FFN (Tancik et al., 2020) and SIREN (Sitzmann et al., 2020), both of which have demonstrated effectiveness across various data types. In our experiments, we apply SIREN for spatially sparse ST and FFN for denser spatial distributions.
Following conventional INR approaches, we assume Gaussian-distributed error in the predicted $\hat{\mathbf{z}}$ and use element-wise mean MSE loss for optimization:
$\mathcal{L}_{\text{embd}}=\frac{1}{|\mathbf{M}_{\mathbf{z}}|}\sum_{\mathbf{M}_{\mathbf{z}}}(\hat{\mathbf{z}}-\mathbf{z}_{\text{gt}})^{2}.$ (2)
While $\hat{\mathbf{z}}$ can serve as the mapping output for downstream analysis, we consider it as an intermediate result, decoding this latent code back into the raw space to achieve a more comprehensive reconstruction.